1, A cylinder with a movable piston contains 255 cm3 air at 100C. At what temperature would you expect the volume to be 155 cm3 if the pressure is kept constant?



1
Expert's answer
2022-04-29T14:36:04-0400

V1 = 255 cm3

T1 = 10°C + 273 = 283 K

P1 = P2 = constant

V2 = 155 cm3

T2 = ???


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the temperature, solve the equation for T2:

T2 = V2T1 / V1

T2 = (155 cm3 × 283 K) / (255 cm3) = 172 K

T2 = 172 K or −101°C


Answer: at 172 K